Can you please show me how to set up this problem so I can solve it:
How much water should be added to a 150ml of a 25% salt solution to make a 15% salt solution?

Answer by stanbon(75887)

How much water should be added to a 150ml of a 25% salt solution to make a 15% salt solution?
---
Equation:
water + water = water
0.75*150 + x = 0.85(150+x)
-----
Multiply thru by 100 to get:
75*150 + 100x = 85*150 + 85x
15x = 10*150
x = 10*10 = 100 ml (amt. of water that must be added to the mixture)
=====================================
Cheers,
Stan H.